# Meta-LoRA: Meta-Learning LoRA Components for Domain-Aware ID Personalization

### v2

@@ 1 Introduction
Text-to-image generative models have seen significant advancements in recent years, demonstrating a remarkable ability to generate high-quality images from textual prompt encodings [22]. Among them, latent diffusion models (LDMs) [23, 28, 15, 21] have proven to be particularly effective, utilizing deep learning to iteratively refine images within a latent space. However, achieving identity personalization -generating images that accurately capture a specific subject’s likeness while maintaining generalization- from a single or few image(s) remains a significant challenge [36].
Figure 1: Our overall training and inference pipeline. In the first stage (meta-learning), we learn a subset of LoRA components capturing domain-specific priors, shared across the identities. In the second stage (personalization), only the identity-specific LoRA components are fit from a single image.
@@
The mainstream approaches to generative model personalization fall broadly into two extremes. On the one end, methods such as DreamBooth [24] and textual inversion [6] rely on general-purpose fine-tuning algorithms for personalization purposes. While effective in many scenarios, these methods rely purely on the priors embedded into the pre-trained generative model. Such priors, however, often provide insufficient specialization for capturing subtle domain-specific nuances, such as fine facial details. To overcome such limitations, it is commonly necessary to use a rich set of subject examples and update significant portions of model parameters in order to achieve accurate identity adaptation. Parameter-efficient fine-tuning techniques, e.g. Low-Rank Adaptation (LoRA) [10] and its variants [8, 1, 16, 35, 12], attempt to reduce adaptation complexity by constraining parameter updates through carefully designed structures; however, since these are general fine-tuning techniques, they share many of the limitations of the traditional fine-tuning techniques.
@@
At the other extreme, feed-forward conditioning-based approaches [33, 17, 20, 9, 30, 31] train ControlNet-like  [34] mechanisms to adapt generative models on-the-fly, circumventing the need for iterative fine-tuning at test time. Although appealing due to their tuning-free adaptation ability, these methods require large-scale training datasets and complex conditioning networks. In addition, their purely feed-forward structure practically limits their ability to capture fine-grained identity details.
@@
In our work, we aim to combine the advantages of these two paradigms, blending the capability of learning domain-specific personalization priors without relying solely on computationally demanding feed-forward conditioning modules. To our knowledge, the only work that closely aligns with this direction is HyperDreamBooth [25]. HyperDreamBooth introduces a hyper-network architecture trained to predict identity-specific LoRA weights from input images, which can then be fine-tuned. However, this design necessitates a large meta-model to map input images to all per-layer LoRA components, inheriting similar drawbacks to pure feed-forward conditioning models: susceptibility to overfitting and challenging training requirements.
@@
To overcome the limitations discussed above, we propose Meta-Low-Rank Adaptation (Meta-LoRA), a novel framework that significantly enhances LoRA-based identity personalization in text-to-image generative models, while maintaining simplicity. Specifically, we propose a structured meta-learning strategy to pre-train a subset of LoRA components in a subject-independent manner, effectively encoding domain-specific priors. This results in a compact, ready-to-adapt LoRA structure that facilitates efficient and precise adaptation to novel identities; avoiding high-complexity modules transforming images to adaptation layers used in prior work [33, 17, 20, 9, 30]. Since our meta-learning yields low-complexity LoRA components, it substantially reduces the required training dataset size compared to prior conditioning-based approaches. Furthermore, the final personalization stage involves a straightforward fine-tuning mechanism that naturally supports a variety of customizations, such as augmentations and using additional losses.
@@
We evaluate the proposed approach with a focus on the FLUX.1-dev model [15], a state-of-the-art text-to-image diffusion system known for its high image quality and strong prompt adherence. Following prior studies, we assess personalized models trained with Meta-LoRA based on (i) the retention of the base model’s original capabilities and (ii) the fidelity of generated faces to target identities. However, our analysis highlights inconsistencies in recent literature regarding dataset choices and evaluation protocols, particularly the problematic reuse of the same image for both personalization and evaluation, which inflates performance estimates. To address these issues, we introduce Meta-PHD, a robust benchmark featuring diverse identities across multiple sources, with varied poses, lighting, and backgrounds, ensuring a rigorous evaluation of identity personalization methods. Our qualitative analysis further complements the quantitative results, capturing subtle yet significant details that conventional metrics may overlook. Together, our results demonstrate that the proposed approach not only preserves identity across diverse scenarios but also maintains strong generalization and prompt adherence, highlighting the potential of meta-learning for enhancing LoRA-based identity personalization.
In summary, our contributions are: (1) a novel Meta-LoRA architecture that separates generic and identity-specific components, (2) a meta-learning strategy, and an effective optimization algorithm, enabling one-shot personalization, (3) a new benchmark dataset (Meta-PHD) for rigorous evaluation, and (4) a Robust Face Similarity metric to better measure identity retention.
@@ 2 Related Works
Identity personalization methods can be grouped into two: fine-tuning and feed-forward conditioning based approaches. In a pioneering work, Ruiz et al. [24] introduces DreamBooth, a fine-tuning approach that balances training with a class-specific prior-preservation loss to maintain model generalization. Custom Diffusion [13] fine-tunes select cross-attention parameters for integrating novel concepts. LoRA [10] offers a parameter-efficient alternative to fine-tuning parameters directly, enabling controlled adaptation by introducing low-rank updates to the model weights. Given its widespread adoption, we focus on the original LoRA formulation, though Meta-LoRA can naturally extend to other LoRA variants, e.g. [1, 8, 16, 35, 12].
@@
In the second group of methods, several feed-forward conditioning methods have recently been proposed (see [36] for a recent survey). For example, PortraitBooth [20], PhotoMaker [17], InstantBooth [27], PhotoVerse [3], FastComposer [32], InstantID [30], and PuLID [9] all incorporate an IP-Adapter-like design, leveraging identity-related embeddings extracted from input ID images to manipulate text-conditioned generation with various enhancements: PortraitBooth and PhotoMaker fine-tune the diffusion model, with PortraitBooth further introducing an identity loss and locational cross-attention control. InstantBooth integrates an adapter layer positioned after the cross-attention mechanism to inject conditional features into U-Net layers. InstantID employs a ControlNet module that incorporates facial landmarks and identity embeddings. FastComposer incorporates subject embeddings extracted from an image encoder in order to augment the generic text conditioning in diffusion models. In PhotoVerse, a dual-branch conditioning mechanism operates across both text and image domains, while the training is further reinforced by a facial identity loss. PuLID enhances identity preservation through ID loss and improves text-image alignment via an alignment loss. In contrast to these works, Meta-LoRA does not require training a complex conditioning model. In addition, we simply use only diffusion loss; although the incorporation of additional losses is an orthogonal direction, and can easily be incorporated into the Meta-LoRA training, if desired.
@@
As a work mixing elements from both groups of personalization approaches, HyperDreamBooth [25] adopts a two-stage pipeline: first, a meta-module is trained to generate LoRA-like weights based on an identity image; second, these generated weights undergo fine-tuning for the specific identity. While this method aligns with our work to some extent, it introduces several limitations that hinder its scalability and applicability; as it requires training of a complex image-to-LoRAs mapping network, expressed by a recurrent transformer architecture. In contrast, we propose to simply meta-learn a subset of LoRA components to embed domain-specific priors.
@@
Finally, we should note that our work has been inspired by the prior work on meta-learning for classification and reinforcement learning tasks. We similarly embrace the meta-learning for fast-adaptation idea pioneered by MAML [5]. The formulation and technical details of Meta-LoRA, however, are vastly different virtually in all major ways.
@@ 3 Methodology
3.1 Preliminaries on Text-to-Image LoRA Training
In diffusion-based text-to-image (T2I) models, Low-Rank Adaptation (LoRA) [10] is primarily integrated into the attention mechanisms of the U-Net backbone, particularly within the cross-attention layers that facilitate interactions between latent image representations and text embeddings. By incorporating low-rank decomposition matrices into the weight update process, LoRA enables efficient fine-tuning while minimizing modifications to the original model parameters. In more advanced architectures, such as Stable Diffusion XL [21] and FLUX.1 [15], LoRA may also be applied to transformer-based components, enhancing latent feature extraction and improving T2I coherence. By limiting adaptation to these critical layers, LoRA optimizes parameter efficiency while maintaining the integrity of the underlying diffusion model.
@@ 3.1 Preliminaries on Text-to-Image LoRA Training
3.2 Meta-LoRA Architecture
We propose a three-layer adaptation framework by re-defining the original LoRA Down-block as an identity-shared adaptation component, which we call LoRA Meta-Down (LoMD), and decomposing the conventional LoRA Up-block into two sub-components, which we call LoRA Mid (LoM) and LoRA Up (LoU). Specifically, given some pre-trained linear layer $W_{0}:\mathcal{R}^{d_{1}}\rightarrow\mathcal{R}^{d_{2}}$ and its input $x$, Meta-LoRA applies the following residual update:
$h=$ $W_{0}x+\Delta Wx$ (1)
$=$ $W_{0}x+L_{\text{up}}^{i}L_{\text{mid}}^{i}L_{\text{meta-down}}x$ (2)
@@ 3.2 Meta-LoRA Architecture
3.3 Training
Our training pipeline comprises two main stages: meta-training and test-time ID personalization. During meta-training, we learn a shared LoRA Meta-Down (LoMD) module for general features while employing distinct LoRA Mid (LoM) and LoRA Up (LoU) layers for character-specific details. In the test-time personalization stage, we fine-tune the LoM and LoU layers based on the given identity image. The following subsections provide a detailed explanation of each stage.
3.3.1 Stage-1: Meta-Training
The key challenge in the meta-training stage is learning a robust meta-representation in LoMD, while preventing under- or overfitting in the identity-specific LoM and LoU layers. A simple training approach could be optimizing all LoRA components simultaneously, for all characters. However, it is not possible to construct batches large enough to cover all identities. Trying to naively remedy this problem by applying stochastic gradient descent to the LoM and LoU components of only the identities available in each batch (and the shared LoMD parameters) also yields poor training efficiency, since the identity-specific parameters are updated much less frequently than the shared ones, resulting in poor gradients for LoMD updates.
@@ 3.3.1 Stage-1: Meta-Training
3.3.2 Stage-2: Test-time ID Personalization
In the second stage, we train only the LoM and LoU layers using a novel target identity image, while keeping the LoMD weights frozen. The $r_{1}$ dimension (i.e., the LoMD output dimension) remains unchanged, whereas $r_{2}$ (i.e., the LoM output dimension) is set to 1 to mitigate overfitting due to the single-image input.
@@ 3.3.2 Stage-2: Test-time ID Personalization
To further prevent overfitting on a single image, we apply data augmentations. First, we detect the face in the image and determine the longer side length of the facial bounding box ($f_{long}$). Then, we create multiple crops of the image using common aspect ratios: 16:9, 4:3, 1:1, 3:4, and 9:16. For each aspect ratio, we center the face and set the shorter side of the cropped image to one of the following values: $f_{long}\cdot$1.5, 2, 2.5, 3.5, 4.5}. If the original image does not support a specific crop size, we select the closest possible size if an image with that size is not extracted. Additionally, we apply random horizontal flips to each image. In total, we generate at most 25 augmented images from a single input.
@@ 4 Experiments
In this section, we evaluate the performance of our approach both quantitatively and qualitatively in comparison to state-of-the-art text-to-image personalization methods. We first introduce our evaluation dataset, experimental setup, and metrics, then present comparative results and discussions.
@@ 4.1 Benchmark data set
To rigorously assess personalization performance, we curated a new test dataset, termed the Meta-LoRA Personalization of Humans Dataset (Meta-PHD). This dataset comprises two complementary parts—drawn from FFHQ [11] and Unsplash-50 [7]—and includes both image and prompt collections designed to challenge the model’s ability to preserve identity and adhere to diverse textual prompts. We will make the Meta-PHD dataset publicly available to facilitate fair comparisons.
Meta-PHD-FFHQ images. The Meta-PHD-FFHQ component contains 60 high-quality face images (30 male and 30 female) selected from the FFHQ test set [11]. These images cover a diverse range of ages and skin tones to ensure broad applicability. We applied strict selection criteria to each image: only one individual is present, the face is unobstructed (i.e., no sunglasses, hats, or hands), the eyes are open, and the head is oriented within 30 degrees of frontal.
Meta-PHD-Unsplash images. This component is inspired by the Unsplash-50 dataset [7] and is sourced from https://unsplash.com/ under permissive licenses. It comprises 98 images collected from 16 identities (with 4–10 images per identity). Within each identity, subjects appear in consistent attire and backgrounds but exhibit varied poses. This design ensures the dataset captures a range of views, thus preventing models from succeeding by merely replicating a single reference pose.
Prompts. For the FFHQ component, we prepared 20 prompts per gender (with some overlap) to test the model’s ability to generate both stylistic transformations and recontextualizations. These prompts, inspired by prior work [25, 30, 33, 9, 27], evaluate both prompt adherence and identity preservation. For the Unsplash component, we designed 10 prompts per identity focusing on pose and background variations while enforcing a frontal, unobstructed view of the subject. An example prompt is: ”A man/woman standing in a sunlit urban plaza, facing directly toward the camera with a fully detailed face and direct eye contact”. The complete prompt list is shared in the Supplementary Material.
Model Base Model Train Dataset CLIP-T (%) $\uparrow$ CLIP-I (%) $\uparrow$ DINO (%) $\uparrow$ R-FaceSim (%) $\uparrow$
InstantID [30] SD-XL [21] 60M 29.37 65.17 53.24 72.09
SD-XL Lightning [18] 30.52 70.79 60.34 75.26
RealVisXL V4 [26] 30.53 70.07 65.17 73.77
PhotoMaker [17] SD-XL 112K 31.51 73.55 68.35 67.57
SD-XL Lightning 31.92 78.28 76.82 63.51
RealVisXL V4 31.64 74.14 77.58 69.14
PuLID [9] SD-XL 1.5M 31.12 69.83 63.53 71.06
SD-XL Lightning 31.86 77.91 73.53 73.50
RealVisXL V4 31.12 73.79 74.67 75.19
FLUX.1-dev [15] 30.95 74.66 74.01 75.72
Rank-1 LoRA 375 Iters [10] FLUX.1-dev 0 31.95 81.96 80.01 57.62
+ augmented inputs FLUX.1-dev 31.88 80.42 80.25 72.69
Rank-1 LoRA 625 Iters FLUX.1-dev 31.63 79.48 77.81 61.41
+ augmented inputs FLUX.1-dev 31.47 77.21 77.33 76.15
Meta-LoRA (ours) FLUX.1-dev 8K (male), 21K (female) 31.66 77.96 77.55 77.16
Table 1: Quantitative comparison of our model with the state-of-the-art publicly available studies in the literature. The scores are taken for five seeds and for both male and female subsets, and averaged to reach the final output. While CLIP-T, CLIP-I and DINO are calculated from the Meta-PHD-FFHQ subset, R-FaceSim is computed using the Meta-PHD-Unsplash images.

### v3

@@ 1 Introduction
Text-to-image generative models have seen significant advancements in recent years, demonstrating a remarkable ability to generate high-quality images from textual prompt encodings clip_paper . Among them, latent diffusion models (LDMs) sd_paper ; sd_35_model ; flux_model ; sdxl_paper have proven to be particularly effective, utilizing deep learning to iteratively refine images within a latent space. However, achieving identity personalization -generating images that accurately capture a specific subject’s likeness while maintaining generalization- from a single or few image(s) remains a significant challenge zhang2024survey .
@@
The mainstream approaches to generative model personalization fall broadly into two extremes. On the one end, methods such as DreamBooth dreambooth_paper and textual inversion gal2022textualinversion rely on general-purpose fine-tuning algorithms for personalization purposes. While effective in many scenarios, these methods rely purely on the priors embedded into the pre-trained generative model. Such priors, however, often provide insufficient specialization for capturing subtle domain-specific nuances, such as fine facial details. To overcome such limitations, it is commonly necessary to use a rich set of subject examples and update significant portions of model parameters in order to achieve accurate identity adaptation. Parameter-efficient fine-tuning techniques, e.g. Low-Rank Adaptation (LoRA) lora_paper and its variants gao24fourierft ; borse2024foura ; yang24vblora ; zhang2023adalora ; kopiczkovera24veralora , attempt to reduce adaptation complexity by constraining parameter updates through carefully designed structures; however, since these are general fine-tuning techniques, they share many of the limitations of the traditional fine-tuning techniques.
Figure 1: Our overall pipeline. In the first stage (meta-learning), we learn a subset of LoRA components capturing domain-specific priors, shared across the identities. In the second stage (personalization), only the identity-specific LoRA components are fit from a single image.
@@
At the other extreme, feed-forward conditioning-based approaches ipadapter_paper ; photomaker_paper ; portraitbooth_paper ; pulid_paper ; instantid_paper ; wei2023elite train ControlNet-like  controlnet_paper mechanisms to adapt generative models on-the-fly, circumventing the need for iterative fine-tuning at test time. Although appealing due to their tuning-free adaptation ability, these methods require large-scale training datasets and complex conditioning networks. In addition, their purely feed-forward structure practically limits their ability to capture fine-grained identity details.
@@
In our work, we aim to combine the advantages of these two paradigms, blending the capability of learning domain-specific personalization priors without relying solely on computationally demanding feed-forward conditioning modules. To our knowledge, the only work that closely aligns with this direction is HyperDreamBooth hyperdreambooth_paper . HyperDreamBooth introduces a hyper-network architecture trained to predict identity-specific LoRA weights from input images, which can then be fine-tuned. However, this design necessitates a large meta-model to map input images to all per-layer LoRA components, inheriting similar drawbacks to pure feed-forward conditioning models: susceptibility to overfitting and challenging training requirements.
@@
To overcome the limitations discussed above, we propose Meta-Low-Rank Adaptation (Meta-LoRA), a novel framework that significantly enhances LoRA-based identity personalization in text-to-image generative models, while maintaining simplicity. Specifically, we propose a structured meta-learning strategy to pre-train a subset of LoRA components in a subject-independent manner, effectively encoding domain-specific priors. This results in a compact, ready-to-adapt LoRA structure that facilitates efficient and precise adaptation to novel identities; avoiding high-complexity modules transforming images to adaptation layers used in prior work ipadapter_paper ; photomaker_paper ; portraitbooth_paper ; pulid_paper ; instantid_paper . Since our meta-learning yields low-complexity LoRA components, it substantially reduces the required training dataset size compared to prior conditioning-based approaches. Furthermore, the final personalization stage involves a straightforward fine-tuning mechanism that naturally supports a variety of customizations, such as augmentations and using additional losses.
@@
We evaluate the proposed approach with a focus on the FLUX.1-dev model flux_model , a state-of-the-art text-to-image diffusion system known for its high image quality and strong prompt adherence. Following prior studies, we assess personalized models trained with Meta-LoRA based on (i) the fidelity of generated faces to target identities and (ii) the prompt following abilities. However, our analysis highlights inconsistencies in recent literature regarding dataset choices and evaluation protocols, particularly the problematic reuse of the same image for both personalization and evaluation, which inflates performance estimates. To address these issues, we introduce Meta-PHD, a robust benchmark featuring diverse identities across multiple sources, with varied poses, lighting, and backgrounds, ensuring a rigorous evaluation of identity personalization methods. Our qualitative analysis further complements the quantitative results, capturing subtle yet significant details that conventional metrics may overlook. Together, our results demonstrate that the proposed approach not only preserves identity across diverse scenarios but also maintains strong generalization and prompt adherence, highlighting the potential of meta-learning for enhancing LoRA-based identity personalization.
In summary, our contributions are: (1) a novel Meta-LoRA architecture that separates generic and identity-specific components, (2) a meta-learning strategy, and an effective optimization algorithm, which enables one-shot personalization and reduces the adaptation time, (3) a new benchmark dataset (Meta-PHD) for rigorous evaluation, and (4) the proposed R-FaceSim metric to measure identity retention in a reliable manner.
@@ 2 Related Works
Identity personalization methods can be grouped into two: fine-tuning and feed-forward conditioning based approaches. In a pioneering work, dreambooth_paper introduces DreamBooth, a fine-tuning approach that balances training with a class-specific prior-preservation loss to maintain model generalization. Custom Diffusion kumari2023multi fine-tunes select cross-attention parameters for integrating novel concepts. LoRA lora_paper offers a parameter-efficient alternative to fine-tuning parameters directly, enabling controlled adaptation by introducing low-rank updates to the model weights. Given its widespread adoption, we focus on the original LoRA formulation, though Meta-LoRA can naturally extend to other LoRA variants, e.g.  borse2024foura ; gao24fourierft ; yang24vblora ; zhang2023adalora ; kopiczkovera24veralora .
@@
In the second group of methods, several feed-forward conditioning methods have recently been proposed (see zhang2024survey for a recent survey). For example, PortraitBooth portraitbooth_paper , PhotoMaker photomaker_paper , InstantBooth instantbooth_paper , PhotoVerse photoverse_paper , FastComposer fastcomposer_paper , InstantID instantid_paper , and PuLID pulid_paper all incorporate an IP-Adapter-like design, leveraging identity-related embeddings extracted from input ID images to manipulate text-conditioned generation with various enhancements: PortraitBooth and PhotoMaker fine-tune the diffusion model, with PortraitBooth further introducing an identity loss and locational cross-attention control. InstantBooth integrates an adapter layer positioned after the cross-attention mechanism to inject conditional features into U-Net layers. InstantID employs a ControlNet module that incorporates facial landmarks and identity embeddings. FastComposer incorporates subject embeddings extracted from an image encoder in order to augment the generic text conditioning in diffusion models. In PhotoVerse, a dual-branch conditioning mechanism operates across both text and image domains, while the training is further reinforced by a facial identity loss. PuLID enhances identity preservation through ID loss and improves text-image alignment via an alignment loss. In contrast to these works, Meta-LoRA does not require training a complex conditioning model. In addition, we simply use only diffusion loss; although the incorporation of additional losses is an orthogonal direction, and can easily be incorporated into the Meta-LoRA training, if desired.
@@
As a work mixing elements from both groups of personalization approaches, HyperDreamBooth hyperdreambooth_paper adopts a two-stage pipeline: first, a meta-module is trained to generate LoRA-like weights based on an identity image; second, these generated weights undergo fine-tuning for the specific identity. While this method aligns with our work to some extent, it introduces several limitations that hinder its scalability and applicability; as it requires training of a complex image-to-LoRAs mapping network, expressed by a recurrent transformer architecture. In contrast, we propose to simply meta-learn a subset of LoRA components to embed domain-specific priors.
@@
Finally, we should note that our work has been inspired by the prior work on meta-learning for classification and reinforcement learning tasks. We similarly embrace the meta-learning for fast-adaptation idea pioneered by MAML maml2017finn . The formulation and technical details of Meta-LoRA, however, are vastly different virtually in all major ways.
@@ 3 Methodology
Preliminaries. In diffusion-based text-to-image (T2I) models, Low-Rank Adaptation (LoRA) lora_paper is primarily integrated into the attention mechanisms of the U-Net backbone, particularly within the cross-attention layers that facilitate interactions between latent image representations and text embeddings. By incorporating low-rank decomposition matrices into the weight update process, LoRA enables efficient fine-tuning while minimizing modifications to the original model parameters. In more advanced architectures, such as Stable Diffusion XL sdxl_paper and FLUX.1 flux_model , LoRA may also be applied to transformer-based components, enhancing latent feature extraction and improving T2I coherence. By limiting adaptation to these critical layers, LoRA optimizes parameter efficiency while maintaining the integrity of the underlying diffusion model.
@@
Meta-LoRA Architecture. We propose a three-layer adaptation framework by re-defining the original LoRA Down-block as an identity-shared adaptation component, which we call LoRA Meta-Down (LoMD), and decomposing the conventional LoRA Up-block into two sub-components, which we call LoRA Mid (LoM) and LoRA Up (LoU). Specifically, given some pre-trained linear layer $W_{0}:\mathcal{R}^{d_{1}}\rightarrow\mathcal{R}^{d_{2}}$ and its input $x$, Meta-LoRA applies the following residual update:
$h=W_{0}x+\Delta Wx=W_{0}x+L_{\text{up}}^{i}L_{\text{mid}}^{i}L_{\text{meta-down}}x$ (1)
@@
Training overview. Our training pipeline comprises two main stages: meta-training and test-time ID personalization. During meta-training, we learn a shared LoRA Meta-Down (LoMD) module for general features while employing distinct LoRA Mid (LoM) and LoRA Up (LoU) layers for character-specific details. In the test-time personalization stage, we fine-tune the LoM and LoU layers based on the given identity image. The following subsections provide a detailed explanation of each stage.
Stage-1: Meta-Training. The key challenge in the meta-training stage is learning a robust meta-representation in LoMD, while preventing under- or overfitting in the identity-specific LoM and LoU layers. A simple training approach could be optimizing all LoRA components simultaneously, for all characters. However, it is not possible to construct batches large enough to cover all identities. Trying to naively remedy this problem by applying stochastic gradient descent to the LoM and LoU components of only the identities available in each batch (and the shared LoMD parameters) also yields poor training efficiency, since the identity-specific parameters are updated much less frequently than the shared ones, resulting in poor gradients for LoMD updates.
@@
Stage-2: Test-time ID Personalization. In the second stage, we train only the LoM and LoU layers using a novel target identity image, while keeping the LoMD weights frozen. The $r_{1}$ dimension (i.e., the LoMD output dimension) remains unchanged, whereas $r_{2}$ (i.e., the LoM output dimension) is set to empirically determined small values to mitigate overfitting due to the single-image input. To further prevent overfitting on a single image, we apply data augmentations. First, we detect the face in the image and determine the longer side length of the facial bounding box ($f_{long}$). Then, we create multiple crops of the image using common aspect ratios: 16:9, 4:3, 1:1, 3:4, and 9:16. For each aspect ratio, we center the face and set the shorter side of the cropped image to one of the following values: $f_{long}\cdot$1.5, 2, 2.5, 3.5, 4.5}. If the original image does not support a specific crop size, we select the closest possible size if an image with that size is not extracted. Additionally, We apply random horizontal flips to each image. In total, we generate up to 25 augmented images from a single input.
@@ 4 Experiments
In this section, we evaluate our approach through both quantitative and qualitative analyses, comparing it against state-of-the-art text-to-image personalization methods. We begin by describing the evaluation dataset, experimental setup, and metrics, followed by a presentation of comparative results and corresponding discussions. The main paper focuses on the ’female’ class, as it provides more instances for stage-1 training. Evaluations on ’male’ are included in the Supplementary Material.
@@ 4.1 Benchmark data set
To rigorously assess personalization performance, we curated a new test dataset, termed the Meta-LoRA Personalization of Humans Dataset (Meta-PHD). This dataset comprises two complementary parts—drawn from FFHQ ffhq_dataset and https://unsplash.com/ and includes both image and prompt collections designed to challenge the model’s ability to preserve identity and adhere to diverse textual prompts. We will make the Meta-PHD dataset publicly available to facilitate fair comparisons.
Meta-PHD-FFHQ images. The Meta-PHD-FFHQ component contains 60 high-quality face images (30 male and 30 female) selected from the FFHQ test set ffhq_dataset . Instead of utilizing the publicly available Unsplash-50 dataset as in the previous studies lcm_lookahead_paper ; pulid_paper , we aim to create a more diverse set that covers different range of ages and skin tones to ensure broad applicability. We applied strict selection criteria to each image: only one individual is present, the face is unobstructed (i.e., no sunglasses, hats, or hands), the eyes are open, and the head is oriented within 30 degrees of frontal.
Meta-PHD-Unsplash images. This component is inspired by the Unsplash-50 dataset lcm_lookahead_paper and is sourced from https://unsplash.com/ under permissive licenses. It comprises 98 images collected from 16 identities (with 4–10 images per identity). Its main novelty is that within each identity, subjects appear in consistent attire and backgrounds but exhibit varied poses. This design ensures the dataset captures a range of views, thus preventing models from succeeding by merely replicating a single reference pose, which is foundational for the proposed R-FaceSim metric.
Prompts. For the FFHQ component, we prepared 20 prompts per gender (with some overlap) to test the model’s ability to generate both stylistic transformations and recontextualizations. These prompts, inspired by prior work hyperdreambooth_paper ; instantid_paper ; ipadapter_paper ; pulid_paper ; instantbooth_paper , evaluate both prompt adherence and identity preservation. For the Unsplash component, we designed 10 prompts per identity focusing on pose and background variations while enforcing a frontal, unobstructed view of the subject. An example prompt is: "A man/woman standing in a sunlit urban plaza, facing directly toward the camera with a fully detailed face and direct eye contact". The complete prompt list is shared in the Supplementary Material.